OpenStudy (anonymous):

using pi, find the area of a circle with a diameter of 12cm

OpenStudy (pinksapphire):

A=πr2

OpenStudy (pinksapphire):

Radius: 6cm

OpenStudy (pinksapphire):

A=π*6*2

OpenStudy (anonymous):

thank you :)

OpenStudy (pinksapphire):

Okay :)

OpenStudy (anonymous):

\[A= \pi \times r ^{2}\]\[r ^{2}=36\]\[\pi \times 36=113.04\]

OpenStudy (anonymous):

pink saphire

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

its pi r squared

OpenStudy (anonymous):

therefore 6 squared is 36 and times pi = 113.04

OpenStudy (anonymous):

@j28319 is correct.

OpenStudy (pinksapphire):

Ops...forgot about that :(

OpenStudy (pinksapphire):

A=πr^2

OpenStudy (anonymous):

@OrangeMaster you too are correct :D We all make mistakes @PinkSapphire human error is natural

OpenStudy (pinksapphire):

A=π*6^2

OpenStudy (pinksapphire):

Thanks for telling me though :)

OpenStudy (anonymous):

thank you! :)

OpenStudy (pinksapphire):

Sorry about that :(

OpenStudy (anonymous):

its cool (^.^)

OpenStudy (anonymous):

wait a minute...

OpenStudy (anonymous):

here are my options: A. A = 36π cm2 B. A = 48π cm2 C. A = 144π cm2 D. A = 6π cm2 ns:

OpenStudy (anonymous):

oh never mind! err mah gerd i was'nt paying attention for a second.:)